40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog, Page 120 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  11958   Thu Jan 28 19:10:16 2016 gautamUpdateLSCStatus of the green PDH circuits

 

Quote:

We will update the X circuit DCC page with an accurate schematic and photo. 

I've uploaded reasonably high-resolution photographs of the uPDH box for the X-end and Y-end on their respective wiki pages. I've uploaded two photos for each box, one of the circuit board (I checked that these photos are clear enough that we can zoom in and read off component values if necessary), and one of the box with the peripherals not integrated into the circuit board (i.e. the minicircuits mixer ZAD-8+ and the little Pomona box that is an LP filter for the output from the mixer). Since I pulled the boxes out, I thought it might not be a bad idea to measure the TFs of these Pomona boxes and make sure nothing weird is going on, I'll put up some plots later. 

Rana and I discussed some things to look at earlier today:

  • Check the voltages at test-points 1,7 and 9, and make sure they are as expected
  • Change the gain of the pre-amp from x2 to x4 - as Eric mentioned in the previous elog, these had already been swapped out. Right now a 600 ohm and 200 ohm resistor pair are being used, so the preamp gain is x4, which should be okay as per the datasheet of the AD8336 VGA amplifier (although the "recommended" resistor pairing is 301 ohm and 100 ohm, but I don't think this is critical?
  • Track down the reason for the difference in Gain settings at the X and Y ends - typically, the X-end PDH box "Gain" knob is set to 10, while that for the Y-end is set to ~4. The fact that the PZT actuator gain for the Y-end is ~5 times larger than for the X-end doesn't seem to account for all of this. As per the attached plot, the difference in gain between the ends is ~35 dB, which is a factor of 50!

I also did a quick check of the behaviour of the Servo Gain potentiometer by checking the resistance at various positions of the knob - we had suspected that the potentiometer may be logarithmic, but I found that it was in fact linear. I'll put up a plot of the gain as a function of the Servo Gain knob position soon,(plot added) along with results from the other checks.

While disassembling the setup at the X-end to get the PDH box out, I noticed that the signal from the LO is going to the mixer through a Pomona box (no such Pomona box is used at the Y-end). I opened it up and found that it contains just a pair of capacitors in parallel, so it's a phase shifter?. The LO signal also goes through an attenuator. The mixer in both boxes is a ZAD-8+, so why is this part of the setup different?

Both PDH boxes are not hooked up at the moment, I will restore the setups at both ends after running a few more checks on the boxes...

Attachment 1: Servo_gain_calibration.pdf
Servo_gain_calibration.pdf
  11957   Thu Jan 28 14:54:49 2016 ericqUpdateLSCStatus of the green PDH circuits

Yesterday, I uploaded some EAGLE schematic files and a LISO source file for the green PDH servo electronics to the 40m LISO git repository. In doing so, I realized that the DCC document for the X box (D1400293) was not updated at the end of the electronics work we did in Aug/Sep 2014. This is entirely my fault. 

The Y box document (D1400294) is currently accurate. 

The missing information is that, as I posted In ELOG 10457, I ended up destroying our original X box, and replaced it with a spare from the ATF. It was restuffed to match the Y end box pretty much exactly. We will update the X circuit DCC page with an accurate schematic and photo. 

Gautam tells me that he and Rana were looking at the outdated schematic and thinking about improvements, but at least some of this was already done back in 2014 (specifically, the resistors used to specify the AD8336 preamp gain were changed).

  11956   Thu Jan 28 00:29:30 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

Summary of the work done today:

Alignment and other work on PSL table

As mentioned in a previous elog, the beatnote amplitude I obtained was tiny - so I checked the alignment of the two beams onto the PD. I did this as follows:

  • Checked the alignment of the two beams on the recombination BS. Moved the steering mirror for the PSL beam until the two were aligned, as verified by eye using an IR card
  • Turn the steering mirror just before the fast focusing lens and thorlabs PD (kept the fork fixed, just loosened the screw on the post to do this) such that the far-field alignment of the two beams could be checked. I used the BS to tweak this alignment as necessary
  • Iterate the previous two steps till I was happy with the alignment
  • Return steering mirror before the PD to its original position, tweak alignment until DC level on the PD was maximized (as verified using an oscilloscope) 
  • Adjust the HWP just after the lightwave laser such that the power arriving at the PD from the PSL beam and the lightwave beam were approximately equal - verified by blocking each beam and checking change in the DC level

After doing all of this, I found a beatnote at ~-10dBm at a temperature of 45.3002 degrees on the Lightwave. The DC level was ~8V (~4V contribution from each beam). 

PLL and frequency nosie measurements:

Pretty much the same procedure as that described in this elog was followed for setting up the PLL and taking the measurements, except that this time, I used the two SR560s in a better way to measure the open loop TF of the PLL. This measurement suggested a UGF of ~ 10kHz, which seems reasonable to me. I turned the 11MHz marconi off because some extra peaks were showing up in the beat signal spectrum. I judged that the beatnote was not large enough to require the use of an attenuator between the PD and the mixer. I was able to lock the PLL easily enough, and I've attached spectra of the control signal (both uncalibrated and calibrated). To calibrate the spectrum, I did a quick check to determine the actuator gain of the spare Lightwave laser, by sweeping the fast PZT with a low frequency (0.5Hz) 1Vpp sine wave, and looking at the peak in the beat signal spectrum move on the network analyzer. This admittedly rough calibration suggests that the coefficient is ~5MHz/V, consistent with the other Lightwave. Eric suggested a more accurate way to do this would be to match up spectra taken using this method and by locking the PLL by actuating on the FM input of the Marconi - I didn't try this, but given the relatively large low-frequency drifts of the beatnote that I was seeing, and that the control signal was regularly hitting ~2V (i.e shifting the frequency by ~10MHz), I don't think this is viable with a low MHz/V coefficient on the Marconi, which we found is desirable as described here

Bottom line:

The spare Lightwave frequency noise seems comparable to the other two measurements (see attachment #2). If anything, it is a factor of a few worse, though this could be due to an error in the calibration? I'm also not sure why the shapes of the spectra from today's measurement differ qualitatively from those in elog 11929 above ~7kHz. 

 

Some random notes:

  • Do we want to do an AM/PM characterization of the spare Lightwave laser as well? It might be easier to do the PM measurement while we have this measurement setup working
  • Yesterday, I noticed some peaks in the spectrum of the PD output while only the PSL beam was incident on it, at ~35MHz and ~70 MHz. They were pretty small (~-50dBm), but still clearly discernible over the analyzer noise floor. It is unclear to me what the source of these peaks are.
Attachment 1: PLL_OLG.pdf
PLL_OLG.pdf
Attachment 2: Freq_noise_comparison.pdf
Freq_noise_comparison.pdf
  11955   Wed Jan 27 23:14:25 2016 gautamUpdatePEMETMX floor vs table noise
Quote:

I didn't really appreciate this measurement until just now. IF you can save the DTT .xml file with all the traces in it (i.e. NOT just the plots), we should save this data for comparison plotting later. Perhaps Gautam can post the gzipped xml file for you into the log.

The accelerometers don't read any real noise below ~3 Hz, so we can't judge the difference down low, but this seems like a good measurement in the 5 - 100 Hz band.yes

Unfortunately I had closed all the DTT windows that Steve had used for the earlier plots. So I took the spectra again - there may be minor differences given that this measurement was taken at ~11pm at night. Anyways, plots and the xml data file are attached.  

 
Attachment 1: X_end_ACC_Data.png
X_end_ACC_Data.png
Attachment 2: X_End_ACC_Data.xml.tar.gz
  11954   Wed Jan 27 20:51:00 2016 ranaUpdatePEMETMX floor vs table noise

I didn't really appreciate this measurement until just now. IF you can save the DTT .xml file with all the traces in it (i.e. NOT just the plots), we should save this data for comparison plotting later. Perhaps Gautam can post the gzipped xml file for you into the log.

The accelerometers don't read any real noise below ~3 Hz, so we can't judge the difference down low, but this seems like a good measurement in the 5 - 100 Hz band.yes

  11953   Wed Jan 27 18:19:45 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...

  11952   Wed Jan 27 08:17:38 2016 SteveUpdatePEMETMX floor vs table noise

 

Quote:

Objective:  measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in

                   the larger optical table

The accelerometrs labeled as MC1 ( just north east  of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:

MC1 to the floor at the north west leg of optical table.

MC2 is in the north east corner of the optical table

Atm2 was taken after table leg bolts were tighed at 40 ft/lb

The spectrum looks similar to ETMY       except  the Z direction 

ETMY .

Conclusion: up to 20 Hz this set up is good.

Attachment 1: ETMX_X_floor_vs_table.png
ETMX_X_floor_vs_table.png
Attachment 2: ETMX_Y_floor_vs_table.png
ETMX_Y_floor_vs_table.png
Attachment 3: ETMX_Z_floor_vs_table.png
ETMX_Z_floor_vs_table.png
Attachment 4: MC1@ETMX.jpg
MC1@ETMX.jpg
Attachment 5: MC2@1Winno.jpg
MC2@1Winno.jpg
Attachment 6: MC2@ETMXtable.jpg
MC2@ETMXtable.jpg
  11951   Tue Jan 26 17:50:22 2016 gautamUpdateGreen LockingLightwave frequency noise measurement

I attempted to measure the frequency noise of the extra Lightwave NPRO we have that is currently sitting on the PSL table. I did the following:

  1. Turn the Lightwave NPRO back on.
  2. Disable MC autolocker and close the PSL shutter.
  3. Checked the alignment of the pick off from the PSL beam and the beam from the Lightwave NPRO onto the PDA10CF. These seemed okay, and I didn't really have to tweak any of the steering optics. I was getting a DC signal level of ~7V (the PD should drive a 1Mohm load up to 10V so it wasn't saturated).
  4. Swept the crystal temperature on the Lightwave using the dial on the front panel of the controller. I found beatnotes at 48.1831 degrees and 45.3002 degrees. However, the amplitude of the beatnote was pretty small (approx. -40dBm on the Agilent NA). I tried playing around with the beam alignment and laser power on the Lightwave NPRO to see if I could increase the beatnote amplitude, but was unsuccessful - turning up the laser power (from the nominal level of 55mW as per the front panel display) caused the PD to saturate at 10V, while as far as I could tell, the alignment of the two beams onto the PD is reasonably good. This seems inconsistent with the numbers Koji has reported in this elog, where he was able to get a beatnote of ~1Vpp for a DC of 2.5 V. 
  5. I tried locking the PLL (in roughly the same configuration as reported here) with this small amplitude beatnote but was unsuccessful. 

I've turned the Lightwave NPRO back to standby for now, in anticipation of further trials later today. I've also restored the IMC. 

  11950   Tue Jan 26 15:07:47 2016 SteveUpdatePEMaccelerometers moved to ETMX

Objective:  measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in

                   the larger optical table

The accelerometrs labeled as MC1 ( just north east  of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:

MC1 to the floor at the north west leg of optical table.

MC2 is in the north east corner of the optical table

Atm2 was taken after table leg bolts were tighed at 40 ft/lb

The spectrum looks similar to ETMY     (Guralp for below 3 Hz )  except  the Z direction 

ETMY  .

Attachment 1: ETMX_floor_vs_optab.png
ETMX_floor_vs_optab.png
Attachment 2: ETMX_floor_vs_optabtop.png
ETMX_floor_vs_optabtop.png
  11949   Mon Jan 25 11:06:08 2016 SteveUpdatePEMnoisy fan belt replaced

The noisy fan belt on the roof of the control room was finally replaced on Friday, Jan 22 2016

Quote:

Air condition maintenance is happening. It should be done by 10am

 

  11948   Mon Jan 25 08:36:46 2016 SteveUpdateSUSPRM damping restored

PRM suspension damping restored after 4.1 Mag Ludlow earthquake.

 

Attachment 1: EQ4.1mLudlow.png
EQ4.1mLudlow.png
Attachment 2: EQ7.1mAlaska_4.1mLudlow.png
EQ7.1mAlaska_4.1mLudlow.png
Attachment 3: 7.1_4.1.png
7.1_4.1.png
  11947   Fri Jan 22 18:46:03 2016 ranaUpdateGreen LockingAUX-X AM/PM investigations

The PDA photodetectors are DC coupled, so you cannot use them to go directly into the analyzer. Must use the DC block so that you can reduce the input attenuation on the B channel and then lower the drive amplitude.

Good policy for TF measurements: drive as softly as you can and still measure in a reasonable amount of time, but no softer than that.

  11946   Fri Jan 22 17:22:06 2016 gautamUpdateGreen LockingAUX-X AM/PM investigations

There were a number of directories in /users/OLD/mott/PZT/2NPRO, I've used the data in Innolight_AM_New. Also, I am unsure as to what their "calibration" factor is to convert the measured data into RIN, so I've just used a value of 0.8, with which I got the plot to match up as close as possible to the plot in this elog. I also redid the measurement today, given that the laser parameters have changed. The main difference was that I used an excitation amplitude of +15dBm, and an "IF Bandwidth" of 30Hz in the parameter files for making these measurements, which I chose to match the parameters Mott used. There does seem to be a shift in some of the features, but the <100kHz area seems similar to the old measurement now. 

Having put the PD back in, I also took measurements of the RIN with the input to the laser PZT terminated. There is no difference with the Noise Eater On or OFF! 

Quote:
Quote:

Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 

It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?

 

Attachment 1: AM_response.pdf
AM_response.pdf
Attachment 2: NE_investigations.pdf
NE_investigations.pdf
  11945   Fri Jan 22 13:33:37 2016 ericqUpdateGreen LockingAUX-X AM/PM investigations
Quote:

Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 

It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?

  11944   Fri Jan 22 11:33:20 2016 gautamUpdateGreen LockingAUX-X AM/PM investigations

I was trying to characterize the AM/PM response of the X end laser. I tried to measure the AM response first, as follows:

  • I used the Thorlabs PDA 55, whose datasheet says it has 10MHz bandwidth - I chose it because it has a larger active area than the PDA 255, but has sufficient bandwidth for this measurement. 
  • My earlier measurement suggested the IR power coming out of the laser is ~300mW. As per the datasheet of the PDA 55, I expect its output to be (1.5 x 10^4 V/A) * (~0.25 A/W) ~ 4000 V/W => I expect the PD output (driving the 50ohm input of the Agilent NA) to saturate at ~1.3mW. So I decided to use a (non-absorptive) ND 3.0 filter in front of the PD (i.e. incident power on the PD ~0.3 mW).
  • I measured the AM response (inputA/inputR) by using the RF output from the Agilent analyzer (divided using a mini-circuits splitter half to input R and half to the laser PZT), and the PD output to input A. I set the power of the RF output on the analyzer to 0 dBm. 
  • Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 
  • I also took a measurement of the RIN with no drive to the laser PZT (terminated with a 50ohm terminator) - see Attachment #2. Qualitatively, this looks like the "free-running" RIN curve on the Innolight datasheet (see Attachment #3, the peak seems slightly shifted to the left though), even though the Noise Eater switch on the laser controller front panel is set to "ON". I neglected taking a spectrum with it OFF, I will update this elog once I do (actually I guess I have to take both spectra again as the laser diode and crystal temperatures have since been changed - this data was taken at T_diode = 28.5deg, I_diode = 1.90A, and T_crystal = 47.5 deg). But does this point to something being broken?
  • I was unable to lock the PLL yesterday to measure the PM response, I will try again today.
Attachment 1: AUX_X_AM.pdf
AUX_X_AM.pdf
Attachment 2: AUX_X_RIN.pdf
AUX_X_RIN.pdf
Attachment 3: NE_Mephisto.png
NE_Mephisto.png
  11943   Fri Jan 22 01:56:01 2016 ericqUpdateLSCAudio ALSX

I hooked up the ALSX DFD output to the fibox, and used the adjustable delay line to set the phase properly. I recorded the noise on pianosa, and have attached it. Of course, this doesn't really capture the low frequency behavior. 

Unrelated to this: I found the MC WFS turned off, and the loops ran away when turning them on. I tweaked the alignment, and reset the WFS offsets. Seems stable for now. 

Attachment 1: ALSX.wav
  11942   Thu Jan 21 18:34:04 2016 ranaUpdateLSCPSL and AUX-X temperatures changed

Is the black ref spectrum from this year or from May of 2015 or ?

I wonder if the noise is a bunch of fast spikes or if its a true broadband rumble. Maybe we can tell by looking at the analog DFD or PLL outputs?

  11941   Thu Jan 21 00:02:11 2016 KojiUpdateLSCHopeful signs

That's a good news. Only quantitative analysis will tell us if it is true or not.

Also we still want to analyze the traffic with the new switch.

Quote:

On a brighter note, I've only noticed one brief EPICS freeze all night. In addition, the wall StripTools seem totally contiuous since ~4pm, whereas I'm used to seeing some blocky shapes particularly in the seismic rainbow. Could this possibly mean that the old WiFi router was somehow involved in all this? 

 

  11940   Wed Jan 20 23:26:10 2016 gautam, ranaUpdateLSCPSL and AUX-X temperatures changed

Earlier today, we did a bunch of stuff to see if we could improve the situation with the excess ALS-X noise. Long story short, here are the parameters that were changed, and their initial and final values:

X-end laser diode temperature:     28.5 degrees ---> 31.3 degrees

X-end laser diode current:             1.900 A ---> 1.942 A

X-end laser crystal temperature:   47.43 degrees ---> 42.6 degrees

PSL crystal temperature:              33.43 degrees ---> 29.41 degrees

PSL Diode A temperature:           21.52 degrees ---> 20.75 degrees

PSL Diode B temperature:           22.04 degrees ---> 21.3 degrees 

The Y-end laser temperature has not yet been adjusted - this will have to be done to find the Y-beatnote.

Unfortunately, this does not seem to have fixed the problem - I was able to find the beatnote, with amplitude on the network analyzer in the control room consistent with what we've been seeing over the last few days, but as is clear from Attachment 1, the problem persists...

 


Details:

  • PSL shutter was closed and FSS servo input was turned off. 
  • As I had mentioned in this elog, the beat can now only be found at 47.41 degress +/- 1 deg, which is a shift of almost 5 degrees from the value set sometime last year, ~42.6 degrees. Rana thought it's not a good idea to keep operating the laser at such a high crystal temperature, so we decided to lower the X-end laser temperature back to 42.6 degrees, and then adjust the PSL temperature appropriately such that we found a beat. The diode temperature was also tweaked (this requires using a small screwdriver to twist the little knob inset to the front panel of the laser controller) - for the end laser, we did not have a dedicated power monitor to optimize the diode temperature by maximizing the current, and so we were just doing this by looking at the beat note amplitude on the network analyzer (which wasn't changing by much). So after playing around a little, Rana decided to leave it at 31.3 degrees.
  • We then went to the PSL table and swept through the temperature till a beat was found. The PMC wouldn't stay locked throughout the sweep, so we first did a coarse scan, and saw weak (due to the PMC being locked to some weird mode) beatnotes at some temperatures. We then went back to 29.41 degrees, and ran the PMC autolocker script to lock the PMC - a nice large beatnote was found. 
  • Finally, Rana tweaked the temperatures of the two diodes on the PSL laser controller - here, the optimization was done more systematically, by looking at the PMC transmitted power on the oscilloscope (and also the MEDM screen) as a function of the diode temperature.
  • I took a quick look at the ALS out of loop noise - and unfortunately, our efforts today does not seem to have noticeably improved anything (although the bump at ~1kHz is no longer there). 

Some details not directly related to this work:​ 

  • There are long cables (routed via cable tray) suitable for RF signals that are running from the vertex to either end-table - these are labelled. We slightly re-routed the one running to the X-end, sending it to the IOO rack via the overhead cable tray so that we could send the beat signal from the frequency counter module to the X-end, where we could look at it using an analyzer while also twiddling laser parameters.
  • A webcam (that also claims to have two-way audio!) has been (re?)installed on the PSL table. The ethernet connection to the webcam currently goes to the network switch on the IOO rack (though it is unlabelled at the moment)
  • The X-end area is due for a clean-up, I will try and do some of this tomorrow. 
Attachment 1: 2016_01_20_ALS_OutOfLoop_1.pdf
2016_01_20_ALS_OutOfLoop_1.pdf
  11939   Wed Jan 20 09:36:11 2016 SteveUpdateSUSETMX damping restored

ETMX suspension damping restored.

 

Attachment 1: EQ.png
EQ.png
  11938   Wed Jan 20 02:53:18 2016 ericqUpdateLSCHopeful signs

[ericq, Gautam]

We gave DRFPMI locking a shot, with the ALS out-of-loop noises as attached. I figured the ALSX noise might be tolerable. 

After the usual alignment pains, we got to DRMI holding while buzzing around resonance. Recall that we have not locked since Koji's repair of the LO levels in the IMC loop, so the proper AO gains are a little up in the air right now. There were hopeful indications of arm powers stabilizing, but we were not able to make it stick yet. This is perhaps consistent with the ALSX noise making things harder, but not neccesarily impossible; we assuredly still want to fix the current situation but perhaps we can still lock.

On a brighter note, I've only noticed one brief EPICS freeze all night. In addition, the wall StripTools seem totally contiuous since ~4pm, whereas I'm used to seeing some blocky shapes particularly in the seismic rainbow. Could this possibly mean that the old WiFi router was somehow involved in all this? 

Attachment 1: 2016-01-20_ALSOOL.pdf
2016-01-20_ALSOOL.pdf
  11937   Tue Jan 19 17:54:39 2016 gautamUpdateLSCALSX Noise still anomalously high

While carrying out my end-table power investigations, I decided to take a quick look at the out-of-loop ALSX noise - see the attached plot. The feature at ~1kHz seems less prominent (factor of 2?) now, though its still present, and the overall noise above a few tens of Hz is still much higher than the reference. The green transmission was maximized to ~0.19 before this spectrum was taken.

EDIT 1130pm: 

We managed to access the trends for the green reflected and transmitted powers from a couple of months back when things were in their nominal state - see Attachment #2 for the situation then. For the X arm, the green reflected power has gone down from ~1300 counts (November 2015) to ~600 counts (january 2016) when locked to the arm and alignment is optimized. The corresponding numbers for the green transmitted powers (PSL + End Laser) are 0.47 (November 2015) and ~0.18 (January 2016). This seems to be a pretty dramatic change over just two months. For the Y-arm, the numbers are: ~3500 counts (Green REFL, Nov 2015), ~3500 counts (Green REFL, Jan 2016) ~1.3 (Green Trans, Nov 2015), ~1 (Green Trans, Jan 2016). So it definitely looks like something has changed dramatically with the X-end setup, while the Y-end seems consistent with what we had a couple of months ago...

Attachment 1: 2016_01_19_ALS_OutOfLoop.pdf
2016_01_19_ALS_OutOfLoop.pdf
Attachment 2: Green_Locking_Trends.png
Green_Locking_Trends.png
  11936   Tue Jan 19 17:27:58 2016 gautamUpdateGreen LockingAUX X power investigations

Last week, Eric and I noticed that the green transmission levels at the PSL table seem much lower now than they did a month or two ago. To investigate this, I attempted to reproduce a power budget for the X endtable setup - see the attached figure (IR powers measured with calorimeter, green powers measured with Ophir power meter). A summary of my observations:

  • The measurements were all made at an AUX-X laser diode current of 1.90A, and laser crystal temperature of 47.41 degrees. The current was chosen on the basis of the AUX-X frequency noise investigations. The temperature was chosen as this is the middle of three end-laser temperatures at wich a beat-note can be found now. Why should this temperature have changed by almost 5 degrees from the value reported here? I checked on the PSL laser controller that the PSL temperature is 33.43 degrees. Turning up the diode current to 2A does not change the situation significantly. Also, on the Innolight datasheet, the tuning geometry graphs' X-axes only runs to 45 degrees. Not sure of what to make of this. I tried looking at the trend of the offset to the slow temperature servo to see if there has been some sort of long-term drift, but was unable to do so...
  • The IR power from the laser seems to have halved, compared to the value in Feb 2014. Is this normal deterioration over two years? Changing the laser diode current to 2A and the laser crystal temperature to ~42 degrees (the conditions under which the Feb 2014 measurements were taken) do not alter these numbers radically.
  • The green power seems to have become 1/4 its value in Feb 2014, which seems to be consistent with the fact that the IR power has halved.

It is worth noting that two years ago, the IR power from the AUX-Y laser was ~280 mW, so we should still be getting "enough" green power for ALS?

 

Attachment 1: X_END_POWER_BUDGET.JPG
X_END_POWER_BUDGET.JPG
  11935   Tue Jan 19 10:25:53 2016 SteveUpdateSUSEQ 3.6M Ludlow

Just an other local earthquake 3.6 Mag Ludlow, Ca

No obvious sign of damage

 

Attachment 1: EQ_3.6M_Ludlow.png
EQ_3.6M_Ludlow.png
  11934   Thu Jan 14 18:41:36 2016 ranaUpdateWienerFilteringNoise Subtraction Puzzler

Just not just pedagogical !  Freq domain MISO coherence based subtraction estimation is much faster than calculating MISO WF. And since each bin is independent of each other, this gives us an estimate of how low the noise can go, whereas the Wiener filter is limited by Kramers-Kronig. We should be able to use this on the L1 DARM channel to do the noise hunting as well as estimating the subtraction efficacy of the pseudo channels that you and Rory come up with.

If you can code up a noise hunter example using DARM + a bunch of aux channels, we could implement it in the summary pages code.

  11933   Thu Jan 14 15:08:37 2016 ericqUpdateWienerFilteringNoise Subtraction Puzzler

The anticlimatic resolution to my subtraction confusion: Spectral leakage around 1Hz. Increasing the FFT length to 256 sec now shows that the FIR WF pretty much achieves the ideal subtraction. 

If nothing else, it's good to have worked out how MISO coherence works.

Attachment 1: subpuzz_resolved.pdf
subpuzz_resolved.pdf
  11932   Thu Jan 14 13:44:05 2016 SteveUpdateSUS earthquake 3.6M

No obvious sign of damage.

 

Attachment 1: 3.6M_JohannesburgCA.png
3.6M_JohannesburgCA.png
Attachment 2: ETMX_3.6_M.png
ETMX_3.6_M.png
  11931   Thu Jan 14 02:33:37 2016 ericqUpdateLSCALSX Noise still anomalously high

[ericq, Gautam]

We checked the UGF of the AUX X PDH servo, found a ~6kHz UGF with ~45 degree phase margin, with the gain dial maxed out at 10.0. Laser current is at 1.90, direct IR output is ~300mW.

We recovered ALS readout of IR-locked arms. While the GTRX seemed low, after touching up the beam alignment, the DFD was reporting a healthy amount of signal. ALSY was perfectly nominal. 

ALSX was a good deal higher than usual. Furthermore, there's a weird shape around ~1kHz that I can't explain at this point. It's present in both the IR and green beats. I don't suspect the DFD electronics, because the Y beat came through fine. The peak has moderate coherence with the AUX X PDH error signal (0.5 or so), but the shape of the PDH error signal is mostly smooth in the band in which the phase tracker output is wonky, but a hint of the bump is present. 

Turning the PDH loop gain down increases the power spectrum of the error signal, obviously, but also smoothens out the phase tracker output. The PDH error signal spectrum in the G=10 case via DTT is drowning in ADC noise a bit, so we grabbed it's spectrum with the SR785 (attachment #2, ASD in V/rtHz), to show the smoothness thereof.

Finally, we took the X PDH box to the Y end to see how ALSY would perform, to see if the box was to blame. Right off the bat, when examining the spectrum of error signal with the X box, we see many large peaks in the tens of kHz, which are not present at the same gain with the Y PDH box. Some opamp oscillation shenanigans may be afoot... BUUUUUT: when swapping the Y PDH box into the X PDH setup, the ~1kHz bump is identical. ugh

Attachment 1: 2016-01-14_ALSXspectra.pdf
2016-01-14_ALSXspectra.pdf
Attachment 2: PDHsig.pdf
PDHsig.pdf
  11930   Wed Jan 13 18:36:00 2016 gautamUpdateLSCrestoration of green beat electronics

In preparation for tonight's work, I did the following:

On the PSL table:

  • Powered the RF amplifiers for the green beat signal on
  • Reconnected the outputs of the Green beat PDs to the RF amplifiers
  • Restored wiring in the fiber box such that both IR beats go to the frequency counter.

At the IOO Rack area:

  • Restored wiring to the frequency counter module such that the IR beats from both arms go to the respective channels
  • Partially cleaned up the setup used for measuring AUX laser frequency noise - moved the SR785 to the X end along with one SR560 so that we can measure the end PDH OLTF
  • Brought the HP network analyzer back to the control room so that we can view the green beatnotes.

At the X-end:

  • Turned the function generator used for PDH locking back on
  • Checked that the AUX laser diode current is 1.90 A, and the crystal temperature is ~47.5 degrees, both of which I think are "good" values from our AUX laser frequency noise measurements
  • Did some minor manual alignment of the PZT mirrors

At the Y-end:

  • Restored the BNC connection from the PDH box to the laser's "FAST" control input. The long BNC cable used for the PLL is still running along the Y-arm, I will clean this up later.

Having done all this, I checked the green transmission levels for both arms (PSL green shutter closed, after running ASS to maximize IR transmission). GTRY is close to what I remember (~0.40) while the best I could get GTRX to is ~0.12 (I seem to remember it being almost double this value - maybe the alignment onto the beat PD has to be improved?). Also, the amplitudes of the beatnotes on the network analyzer are ~-50dBm, and I seem to remember it being more like -25dBm, so maybe the alignment on the PD is the issue? I will investigate further in the evening. It remains to measure the OLTF of the X-end PDH as well.

  11929   Tue Jan 12 19:38:31 2016 gautamUpdateLSCFrequently making noise

EDITS 15Jan:

  1. Schematic of test setup added (Attachment #5). Note that the UGF measurements were made with the LPF and gain on the 'wrong' SR560, in a way defeating the purpose of having 2 SR560s in the setup. I only realised this after taking the measuements. But having done the loop algebra, I believe we can extract the necessary information, which is what has been done in subsequent plots...
  2. Koji pointed out that UGFs of ~100kHz was probably too high - this is when I took a closer look at the setup and realised the remarks made above in point 1. I realised we were in fact measuring the 'Process' open-loop TF. We can recover the loop TF by measuring the controller TF (which I did, see Attachment #3). The UGF for the PSL+X PLL loop is ~7.5kHz while that for PSL+Y is ~22kHz (both with a 1Hz LP on the SR560 and gain of x200).
  3. During the above investigations, I found that the measured TF for a 1Hz LP on the SR560 is weird - there seems to be a zero around 5kHz which gives some phase lead where one would expect a uniformly decaying gain and phase to be -90 degrees. Eric and I confirmed this behavioud on another SR560. Low-pass at 10kHz and high-pass at 1kHz seem to work fine. I will investigate this further when I get the time. Anyhow I don't think this affects anything as long as we measure the correct OLTF. It is still not clear to me why we even need this to lock the PLL...
  4. All the spectra (Attachment #4 and #5) are now calibrated taking into account the loop TF. I've added another panel with the spectra in V/rtHz as measured on the SR785, along with the SR560 output noise. I don't think any of the conclusions below are affected by these edits.

Summary:

I took several measurements today using the revised PLL scheme of using the Marconi just as an LO, and actuating on the Laser PZT to keep the PLL locked (I will put up a sketch soon). On the evidence of the attached plots (spectra of PLL control signal), I guess we can conclude the following:

  1. The AUX X laser's frequency noise performance is consistent with the levels expected from 'typical' NPRO numbers (and the datasheet), and is more or less consistent across different diode currents/crystal temperatures (? see below...).
  2. The diode current should be set to something less than 2.00 A
  3. Qualitatively, there is a difference in the shape of the spectra between the PSL+X and PSL+Y combinations above a couple of kHz. I don't know why we see this.

Attachment #2: Measured OLG of PLL for the PSL+X and PSL+Y combinations. The UGF in both cases looks to be above 100 kHz, so I didn't do any calibration for the spectra attached. The gain on the SR560 was set to 200 for all measurements.

Attachment #3: Measured spectra of PLL control signal for various diode currents, with one reading from the PSL+Y combination plotted for comparison. When we took some data last night, Eric noted that there was a factor of ~6 increase in the overall frequency spectrum level at higher currents, I will update the plots with last night's data as well shortly. I found it hardest to keep the PLL locked at a diode current of 2.00 A across all measurements.

Attachment #4: Measured spectra of PLL control signal at two different crystal temperatures. There does not seem to be any significant dependance on temperature, although I did only do the measurement at two temperatures.

Attachment #4 Attachment #1All the data used to make these plots (plus some that have yet to be added to the plots, I will update them).

Misc notes:

  • All measurements taken with two free-running lasers (PSL shutter closed)
  • The SR560 noise was measured with the input on the SR560 set to ground. 
  • In order to go from V/rtHz to Hz/rtHz on the plots, I used 1MHz/V for the X-end laser (which I verified by a quick measurement today to be approximately correct) and 4.6 MHz/V for the Y-end laser, based on an earlier measurement. 
  • I re-routed the long BNC cable to the Y-end, have yet to remove it. The BNC from the PDH setup at the X-end has been re-attached to the X-end NPRO.

Unrelated to this work:

When I came in this afternoon, I noticed that the PMC was unlocked. The usual procedure of turning the servo gain to -10dB and playing around with the DC output adjust slider on the MEDM screen did not work. Eric toggled a few buttons on the MEDM screen after which we were able to relock the PMC using the DC output adjust slider.

Attachment 1: 2016_01_AUXLaser.tar.gz
Attachment 2: OLGs.pdf
OLGs.pdf
Attachment 3: variedCurrent.pdf
variedCurrent.pdf
Attachment 4: variedTemp.pdf
variedTemp.pdf
Attachment 5: PLL_setup.pdf
PLL_setup.pdf
  11928   Tue Jan 12 08:40:06 2016 SteveUpdatePEMAir cond maintenance

Air condition maintenance is happening. It should be done by 10am

  11927   Tue Jan 12 07:54:23 2016 SteveUpdateSUSETMX damping restored

PMC locked and ETMX suspension damping restored.

  11926   Tue Jan 12 03:03:55 2016 ericqUpdateLSCFrequently making noise

Gautam will soon follow up with detailed analysis, but here is a brief summary of some of our activities and findings.

  • Two Marconis were beat together in various ways, we figured the noise added by turning on external modulation didn't make us happy. 
  • I locked the AUX X laser to the PSL via PZT. I'm more likely to believe we're seeing real broadband laser noise in this configuration; locking the the PSL laser to the IMC brought the noise down in a reasonable way. The PLL bandwidth was a smidge over 100k.
  • We saw a factor ~6 increase in noise when changing the diode current from 1.8 to 1.96A. We'll be following this up at more temperatures and currents soon. 
  • Gautam will verify the AUX X laser PZT calibration tomorrow, and post calibrated spectra of this increase. 

Please note that there is a long BNC cable still laid out from the IOO rack area to the X end table; watch your step!

  11925   Mon Jan 11 19:01:56 2016 gautamUpdateLSCPLL Marconi Investigation

EDIT 01/12/2016 6PM: I've updated the plots of the in-loop spectra such that they are calibrated throughout the entire domain now. I did so by inferring the closed-loop transfer function (G/(1-G)) from the measured open-loop transfer function (G), and then fitting the inferred TF using vectfit4 (2 poles). The spectra were calibrated by multiplying the measured spectra by the magnitude of the fitted analytic TF at the frequency of interest.

EricQ brought back one of the Marconis that was borrowed by the Cryo lab to the 40m today (it is a 2023B - the Marconi used for all previous measurements in this thread was 2023A). Koji had suggested investigating the frequency noise injected into the PLL by the Marconi, and I spent some time investigating this today. We tried to mimic the measurement setup used for the earlier measurements as closely as possible. One Marconi was used as a signal source, the other as the LO for the PLL loop. All measurements were done with the carrier on the signal Marconi set to 310MHz (since all our previous measurements were done around this value). We synced the two Marconis by means of the "Frequency Standard" BNC connector on the rear panel (having selected the appropriate In/Out configurations digitally first). Two combinations were investigated - with either Marconi as LO and signal source. For each combination, I adjusted the FM gain on the Marconi (D in the plot legends) and the overall control gain on the SR560 (G in the plot legends) such that their product remained approximately constant. I measured the PLL OLG at each pair to make sure the loop shape was the same throughout all trials. Here are the descriptions of the attached plots:

Attachment #1: 2023A as LO, 2023B as source, measured OLGs

Measured OLG for the various combinations of FM gain and SR560 gain tested. The UGF is approximately 30kHz for all combinations - the exceptions being D 1.6MHz, G=1e4 and D=3.2MHz, G=1e4. I took the latter two measurements just because these end up being the limiting values of D for different carrier frequencies on the Marconi.

Attachment #2: 2023A as LO, 2023B as source, measured spectra of control signal (uncalibrated above 30kHz)

I took the spectra down to 2Hz, in two ranges, and these are the stitched versions. 

Attachment #3: 2023B as LO, 2023A as source, measured OLGs

Attachment #4: 2023B as LO, 2023A as source, measured spectra of control signal (uncalibrated above 30kHz)

So it appears that there is some difference between the two Marconis? Also, if the frequency noise ASD-frequency product is 10^4 for a healthy NPRO, these plots suggest that we should perhaps operate at a lower value of D than the 3.2MHz/V we have been using thus far? 

As a quick trial, I also took quick spectra of the PLL control signals for the PSL+Aux X and PSL+Aux Y beat signals, with the 2023B as the LO (Attachment #5). The other difference is that I have plotted the spectrum down to 1 Hz (they are uncalibrated above 30Hz). The PSL+Y combination actually looks like what I would expect for an NPRO (for example, see page 2 of the datasheet of the Innolight Mephisto) particularly at lower frequencies - not sure what to make of the PSL+X combination. Also, I noticed that the amplitude of the PSL+Y beatnote was going through some large-amplitude (beat-note fluctuates between -8dBm and -20dBm) but low frequency (period ~10mins) oscillations. This has been observed before, not sure why its happening though. 

More investigations to be done later tonight.

Attachment 1: 2023ALockedto2023B.pdf
2023ALockedto2023B.pdf
Attachment 2: 2023ALockedto2023B_spectra.pdf
2023ALockedto2023B_spectra.pdf
Attachment 3: 2023BLockedto2023A.pdf
2023BLockedto2023A.pdf
Attachment 4: 2023BLockedto2023A_spectra.pdf
2023BLockedto2023A_spectra.pdf
Attachment 5: TestSpectra.pdf
TestSpectra.pdf
Attachment 6: 2016_01_AUXLaser.tar.gz
  11924   Sat Jan 9 00:39:15 2016 gautamUpdateLSCAUX Y Freq Noise re-measured
Quote:

With the Y end laser, I was able to lock the PLL with a lower actuation range (1.6MHz/V), and with the PSL in both the free-running and MCL locked configurations.

I took spectra (attached) with the same actuation range (3.2 MHz/V) for the AUX X+PSL and AUX Y+PSL combinations (PSL shutter closed) just to keep things consistent. It looks like there is hardly any difference between the two combinations - could the apparent factor of 3 worse performance of the X end laser have been due to different actuation ranges on the Marconi? 

I've not managed to take a spectrum for the proposed replacement Lightwave laser on the PSL table, though with Eric's help, I've managed to find the beatnote (at a temperature of 53.0195 degrees). I had to do some minor alignment tweaking for this purpose on the PSL table - the only optics I touched were the ones in the pink beam path in attachments 1 and 2 in this elog (the setup used to make the measurement is also qualitatively similar to attachment 3 in the same elog, except for the fact that we are feeding back to the Marconi and not the laser - a detailed sketch with specific components used will be put up later). I'll try and measure the frequency noise of this laser as well over the weekend and put up some spectra. 

With regards to possibly switching out the Lightwave on the PSL table for the (faulty?) Innolight at the X end, I've verified the following:

  • The beam-height from the Lightwave on the mount it is currently sitting on is the same as that from the Innolight on the X end table.
  • There is sufficient space on the X end table to house the Lightwave laser+mount

It remains to characterize the beam coming out from the Lightwave laser and do a mode matching calculation to see if we can use the same optics currently in place (with slight rearrangement) to realize a satisfactory mode-matching solution - I've obtained a beam profiler to do this from Liyuan and have the software setup, but have yet to do the beam scan - the plan is to do this on the SP table, but we've put off moving the Lightwave laser off the PSL table until we (i) establish conclusively that the X end laser is malfunctioning and (ii) check the frequency nosie of the Lightwave relative to the Aux lasers currently at the ends. 

The area around the Marconi is in a little disarray at the moment with a bunch of cables, SR560s, analyzers etc - I didn't want to disconnect the measurement setup till we're done with it. I have however turned both IR beat PDs on the PSL table off, and have reconnected the Marconi output to the Frequency Generation Unit and have set the carrier back to 11.066209MHz, +13dBm. 

Attachment 1: AuxPLL.pdf
AuxPLL.pdf
  11923   Fri Jan 8 22:37:17 2016 KojiUpdateGeneralNew WiFi router

I configured a new wifi bridge for a GPIB Instruments.

The some facts are described on https://wiki-40m.ligo.caltech.edu/Network

The setting up wasn't so straight forward. I added more details there as a linked page.
One thing I had to do with the martian wifi router was that I had to separate the name of SSIDs for 2GHz and 5GHz networks.


Now the data download from Agilent is super fast!
The first establishing the connection takes the most of the time, and the data transfer takes literary nothing.

controls@pianosa|netgpibdata > time ./netgpibdata -i 192.168.113.167 -d AG4395A -a 10 -f meas01
Connecting to host 192.168.113.167, GPIB 10...
done.
Data will be written into meas01.dat.
Parameters will be written into meas01.par.
Writing measurement data to file...
Writing to the parameter file.

real    0m4.056s
user    0m0.068s
sys    0m0.020s

 

  11922   Fri Jan 8 20:02:49 2016 ranaUpdateLSCAUX Y Freq Noise measured

Unless this is the limit from the way you guys set up the PLL, it seems like there's no difference between the two lasers that's of any import. So then the locking problem has been something else all along - perhaps its noise in the X-PDF lock somehow? PDH box oscillations?

  11921   Fri Jan 8 14:47:33 2016 ericqUpdateLSCAUX Y Freq Noise measured

Here are some results from measuring the PSL / AUX Y beat. 

With the Y end laser, I was able to lock the PLL with a lower actuation range (1.6MHz/V), and with the PSL in both the free-running and MCL locked configurations. (In the latter, I had to do a bit of human-turning-knob servo to keep the control signal from running away). I also took a spectrum with the marconi detuned from the beat frequency, to estimate the noise from the PD+mixer+SR560. 

It looks like the AUX X laser is about 3 times noisier than the Y, though the Y laser looks more like a 10^5 noise-frequency product, whereas I thought we needed 10^4. 

Gautam is investigating the PSL / AUX PSL beat with Koji's setup now. 

Attachment 1: AUX_freqnoises.pdf
AUX_freqnoises.pdf
Attachment 2: AUXY_Jan8.zip
  11920   Thu Jan 7 19:04:25 2016 KojiUpdateLSCAUX X Freq Noise measured

The next step is to compare this data with the same measurement with the PSL and the AUX laser on the PSL table (or the end Y laser). If these show a lot lower noise level, we can say 1) the x-end laser is malfunctioning and 2) the y-end and AUX laser on the PSL are well low noise.

  11919   Thu Jan 7 16:52:32 2016 ericqUpdateLSCAUX X Freq Noise measured

Here is some of the promised data. As mentioned, changing diode current and crystal temperature didn't have much effect on the frequency noise spectrum; but the spectrum itself does seem too high for our needs. 

At each temperature, we started measuring the spectrum at 1.8A, and stepped the current up, hoping to reach 2.0 A.

At 47.5 C, we were able to scan the current from 1.8 to 2.0 A without much problem. At 49.0C, the laser mode would hop away above 1.95A. At 50.4C it would hop away above 1.85A. The spectra were not seen to change when physically disconnecting the PZT actuation BNC from the rear of the laser. 

The flattening out at the upper end is likely due to the SR560 output noise. I foolishly neglected to record the output spectrum of it, but with the marconi external modulation set to 3.2MHz/V, the few Hz/rtHz above 20k translates to a signal on the order of uV/rtHz, which seems reasonable. 

Data and code attached. 

Attachment 1: AUXfreqnoise.pdf
AUXfreqnoise.pdf
Attachment 2: auxXmeasurements.zip
  11918   Thu Jan 7 15:29:54 2016 ericqUpdateWienerFilteringNoise Subtraction Puzzler

The puzzle continues...

I found some reference for computing "multicoherence," which should properly estimate the potential MISO subtraction potential in situations where the witness channels themselves have nontrivial coherence. Specifically, I followed the derivations in LIGO-P990002. The underlying math is related to principal component analysis (PCA) or gram-schmidt orthogonalization. 

This produced the following results, wherein the Wiener subtraction is still below what the coherences predict. 

I've attached the data and code that produced this plot. 

Attachment 1: subpuzz2.pdf
subpuzz2.pdf
Attachment 2: puzzle.zip
  11917   Thu Jan 7 04:28:39 2016 ericqUpdateLSCAUX X Freq Noise measured

[ericq, Gautam]

Brief summary of tonights work:

  • Locked Marconi to AUX X vs PSL beat at around 320MHz, PSL shutter closed (i.e. both lasers free running)
  • Measured control signal spectrum at various laser diode currents, crystal temperatures. Oddly, spectra remained consistent across these variables. 
  • Measured OLG of PLL to calibrate into open-loop frequency noise of the beat, found UGF ~30kHz

Our "requirement" for the end laser is as follows: We expect to (and have in the past) achieved ALS sensitivity of 1Hz/rtHz at 100 Hz. If the end PDH loop is 1/f from 100Hz-10kHz, then we have 40dB of supression at 100Hz, meaning the free running AUX laser noise should be no more than 100Hz/rtHz at 100Hz.

So, if we expect both the PSL and AUX lasers to have this performance when free running, we would get the green curve below. We do not. frown


I'll post more details about the exact currents, temperatures and include calibrated plots for the >30kHz range later. Here's the OLG for kicks. 

Attachment 1: PLLspec.pdf
PLLspec.pdf
Attachment 2: PLL_OLG.pdf
PLL_OLG.pdf
  11916   Wed Jan 6 16:40:49 2016 ericqUpdateGeneralNew WiFi router

The new wifi router, a Netgear R6400, has been installed, next to the old one which is disconnected (but not yet removed). 

Same SSID, and I've added only the wireless MAC addresses of viviana, paoloa and asia, the three thinkpads inside.

Qualitatively, dataviewer at the X end seems pretty snappy. I'll do some more quantitative comparison of the two routers at some point soon. I will update the wiki, too.

  11915   Wed Jan 6 10:58:03 2016 SteveUpdateGeneral projector light bulb is out

Projector light bulb ordered

Quote:

 

[ Manasa, Ericq and Steve ]

 Vivitek D952HD with 186 hours installed.

 

  11914   Wed Jan 6 08:09:04 2016 SteveUpdatePEM4.5 Mag Banning EQ

IFO restored after 4.5 Mag Banning, Ca earthquake.

Attachment 1: 4.5Mag__Banning.png
4.5Mag__Banning.png
  11913   Tue Jan 5 17:19:57 2016 ranaUpdatesafetyRat trap locations

In the modern times, people use glue traps to catch rats instead of springs. They are less hazardous to people and don't spread rat fluid on the floor.

  11912   Tue Jan 5 16:33:45 2016 ericqUpdateLSCAUX X Freq Noise attempt

Turning on the MCL path (in addition to the MCL FF we always have on) let me lock the PLL for multiple seconds, but low frequency excursions still break it in the end. I was able to briefly observe a level of ~50Hz/rtHz at 1kHz, which may or may not be real. Tomorrow we'll send the PLL control signal to MC2, which should lock it up just fine and give us time to twiddle laser diode current, measure the PLL loop shape, etc. 

  11911   Tue Jan 5 15:48:16 2016 SteveUpdatesafetyRat trap locations

Please look around when working close to these five locations. Use flashlights or leave lights on.

These mechanial traps are HAZARDOUS !

No visitors or tours till Monday, Jan 11  2016

Attachment 1: Control_Room.jpg
Control_Room.jpg
Attachment 2: EarmCESdoor.jpg
EarmCESdoor.jpg
Attachment 3: 1Y1.jpg
1Y1.jpg
Attachment 4: MC2.jpg
MC2.jpg
Attachment 5: XarmMaglev.jpg
XarmMaglev.jpg
  11910   Tue Jan 5 13:17:06 2016 ranaUpdateLSCAUX X Freq Noise attempt

The problem here is that the MC displacement noise is leading to large frequency excursions of the PSL beam. Options

  1. Feed back the low frequency PLL control signal to the MC2 length to suppress the excursion required by the Marconi. This is better than driving the laser, since the drive to the laser would be squashed by the MC locking loop.
  2. Put the beat signal through a divider? Don't know if this makes the Marconi more able to handle it.
  3. Turn on the MCL path. this will make the low frequency MC error signal go to the MC length, thereby reducing the low frequency feedback to the NPRO.
  11909   Tue Jan 5 09:48:52 2016 SteveUpdatePEMRat trap moved

Our janitor confirmed that Q was not hallucinating about this animal. The dropping size indicating a good size one in the IFO room.

One of the mechanical traps moved from the control room to the east arm, close to the " machine shop " door.

I'm going to get more traps.

Quote:

Two mechanical and two sticdky traps were set to catch univited visitor.

Absolutely no food or food remains into inside garbage cans!!!!!!!!!!!!!!!!!!!!!!!!!

Quote:

A small rat / large mouse just ran through the control room. Ugh.

 

 

ELOG V3.1.3-